The emf of cell Ag|AgSO4 (0.0083M) || AgSO4 (xM)| Ag is 0.074V at 250C. Find the value of X
1
Expert's answer
2020-05-26T14:20:37-0400

Solution.

"E = \frac{0.059}{n} \times lg(\frac{C(cathode)}{C(anode)})"

"0.074 = \frac{0.059}{1} \times lg(\frac{X}{0.0083})"

X = 0.15 M

Answer:

X = 0.15 M